You’re using a winch to lift supplies that weigh 2380 N a distance of 6.80 meters. How much time does it take an 820 Watt motor to lift these supplies?



1
Expert's answer
2021-02-28T07:33:50-0500

The power

"P=\frac{work \; done}{time}=\frac{Fd}{t}"

Hence, the time

"t=\frac{Fd}{P}=\frac{2380*6.80}{820}=19.7\:\rm s"